A jet can travel 4550 km in 5 hours with the wind. The return trip against the wind takes 6.5 hours. Find the rate of the jet with no wind.
Found 2 solutions by lynnlo, lwsshak3:
Answer by lynnlo(4176)

791.3km without wind

Answer by lwsshak3(11628)

A jet can travel 4550 km in 5 hours with the wind. The return trip against the wind takes 6.5 hours. Find the rate of the jet with no wind.
**
let x=rate of jet with no wind
let c=rate of wind
(x+c)=rate of jet with wind
(x-c)=rate of jet against wind
travel time=distance/rate
..
4550/(x+c)=5
4550/(x-c)=6.5
..
5x+5c=4550
6.5x-6.5c=4550
..
32.5x+32.5c=29575
32.5x-32.5c=22750
add
65x=52325
x=805
rate of jet with no wind=805 km/hr